OpenStudy (loser66):

Need help putting the logic in neat. :) Let G, H,K be groups f1: G--> H f2: G--> K f3: G--> H x K f3 sends g to (f1(g) , f2(g)) f1, f2 are homomorphisms Prove f3 is homomorphisms. I have it done then prove kernel f is \(kernel ~f1\cap kernel~f2\) Please, help

OpenStudy (loser66):

It doesn't allow me post the long comment, ha

OpenStudy (anonymous):

Kernel doesn't make sense unless you have more structure. Are they groups?

OpenStudy (loser66):

yes, they are

OpenStudy (anonymous):

You should state that in your question next time.

OpenStudy (loser66):

:) I think people can help me in this site know what I am doing. hihihi... Since not many of helpers here help me

OpenStudy (anonymous):

Do you know that an isomorphism has a trivial kernel?

OpenStudy (loser66):

\[kernel f_1=\{g\in G |~f_1(g) = e_H\}\]

OpenStudy (loser66):

oh, I am so sorry

OpenStudy (anonymous):

Since it is an isomorphism the kernel contains only one element, the identity element of G.

OpenStudy (loser66):

They are not iso, they are just homomorphism

OpenStudy (anonymous):

Can you edit your question to correct that?

OpenStudy (loser66):

\[kernelf_2=\{g\in G| f_2(g) =e_K\}\]

OpenStudy (anonymous):

In your question is states: "f1, f2 are isomorphisms,"

OpenStudy (loser66):

I did

OpenStudy (loser66):

This stupid computer limits my typing

OpenStudy (loser66):

\[kernel~f_1\cap kernel~f_2 =\{g\in G ~| ~f_1(g)=e_H \cap f_2(g) = e_K\}\]

OpenStudy (anonymous):

So let \(a, b \in G\) we need to show that \(f_3(a \cdot b) = f_3(a)\cdot f_3(b)\). So let us verify. $$\begin{eqnarray*} f_3(a\cdot b) &=& (f_1(a \cdot b), f_2(a \cdot b)) \\ &=& (f_1(a)\cdot f_1(b), f_2(a)\cdot f_2(b)) \\ &=& (f_1(a), f_2(a)) \cdot (f_1(b), f_2(b)) \\ &=& f_3(a)\cdot f_3(b) \end{eqnarray*}$$ So \(f_3\) is a homomorphism.

OpenStudy (loser66):

but \(f_1(g)= e_H=kernel~ f_1\)

OpenStudy (anonymous):

Now we move to the next part.

OpenStudy (loser66):

Yes, I did the same for that part.

OpenStudy (anonymous):

Well the second part is really easy to see. Where are you having trouble?

OpenStudy (loser66):

I want the smoothly logic from kernel f3 to intersection of kernel f1 and kernel f2.

OpenStudy (loser66):

by symbols, not verbal.

OpenStudy (anonymous):

So, we start with the identity element in \(H \times K\) which is \((e_H, e_K)\). Consider the preimage of the identity element.

OpenStudy (loser66):

I mean I don't want to verbalize the logic.

OpenStudy (loser66):

\(f_3^{-1}(e_H,e_K) = kernel f_3\)

OpenStudy (loser66):

\(=\{g\in G~| f(g) =(e_H,e_K)\}\)

OpenStudy (loser66):

I mean f_3

OpenStudy (anonymous):

I don't know how to avoid a bit of English, but it is clear that you end up with the set \(\{g \in G : f_1(g) = e_H \wedge f_2(g) = e_K\}\)

OpenStudy (loser66):

Oh, so we just jump to this?? no need to put anything else between them? Although I don't like to use verbal in Maths, but if it is needed, I can do it. :)

OpenStudy (loser66):

I think I got it. Thanks for your help.